7. A firm is considering replacing an existing distillation column for separating benzene and toluene. One column uses traditPLEASE SHOW ALL WORK NOT USING EXCEL
Answer #1

Assumption: Repeatability of the alternatives for any number of times and that the costs involved would remain the same.

Given,

MARR = 8% per year

Packed Column:

First Cost = -$787,100

Annual Benefits = $78,000

Salvage Value = $99,000

Annual maintenance costs = -$4,000

Estimated life = 45 years

From the compound interest table, we obtain

(A/P, 8%, 45) = 0.0826

(A/F, 8%, 45) = 0.00259

Annual worth of 'packed column' = AW of the first cost + Annual benefits + Annual maintenance cost + AW of ṭhe salvage value

=-$787,100(A/P, 8%, 45) + $78,000 - $4,000 + $99,000(A/F, 8%, 45)

= -$787,100 * 0.0826 + $78,000 - $4,000 + $99,000*0.00259

= $9,241.95

AW of the 'packed column' alternative = $9,241.95

Tray Column:

First Cost = -$684,220

Annual Benefits = $70,000

Salvage Value = $40,000

Annual maintenance costs = -$2,000

Estimated life = 34 years

From the compound interest table, we obtain

(A/P, 8%, 34) = 0.0863

(A/F, 8%, 34) = 0.0063

Annual worth of 'tray column' = AW of the first cost + Annual benefits + Annual maintenance cost + AW of ṭhe salvage value

= -$684,220(A/P, 8%, 34) + $70,000 - $2,000 + $40,000(A/F, 8%, 34)

= -684,220*0.0863 + $70,000 - $2,000 + $40,000*0.0063

= $9,203.814

AW of the 'tray column' alternative = $9,203.81

As the equivalent annual worth of 'Packed column' alternative is higher than that of the 'tray column' alternative, Packed column alternative is recommended.